OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations eliminated today? 430 720 1150 1300

OpenStudy (anonymous):

a + b = 1150 3a - 2b = 1300 2 * (a + b) + 3a - 2b = 2 * 1150 + 1300 2a + 2b + 3a - 2b = 2300 + 1300 5a = 3600 5a/5 = 3600/5 a = 720 a + b = 1150 720 + b = 1150 b = 430

OpenStudy (anonymous):

A. 430